# On long-duration storage, weather uncertainty and limited foresight

### v1

@@ 5.2 Implications
There is broad literature consensus that the design of variable renewable energy systems requires taking into account interannual weather variability [pfenninger_dealing_2017, collins_impacts_2018, ruggles_planning_2024, gotske_designing_2024, goke_stabilized_2024, grochowicz_intersecting_2023, grochowicz_using_2024], especially in the context of long-duration energy storage (LDES) serving as backup during prolonged periods of energy scarcity [brown_ultra-long-duration_2023, dowling_role_2020]. However, the role of limited foresight with respect to weather uncertainty in capacity expansion models is often ignored. Few studies have considered limited foresight in LDES operations and have either done so by using heuristic bidding rules [brown_price_2025, dupre_la_tour_towards_2023], or by using very reduced scenario trees [blanchard_strategic_2024, hummelen_exploring_2024] in partial capacity expansion models.
@@
Ignoring weather uncertainty capacity expansion planning can lead to biased systems designs and operations. Perfect foresight models tend to rely too much on wind capacities as they can perfectly leverage intermittent windy episodes throughout an operational year. In a limited foresight model, solar PV gains in system value as its capacity factors are less variable, aiding the LDES stockpiling process. Meanwhile, LDES capacities themselves do not change much and energy capacities remain close to those defined by the largest dispatch events such as the winters of 1996/97 and 2010/11 under PF as found in previous literature contributions [kittel_coping_2025, ruhnau_storage_2022].
@@
It is obvious that highly-resolved sector-coupled capacity expansion models would be intractable to solve with a stochastic dynamic (dual) programming method like we apply here. Yet, the derived LDES bidding curves may serve as a starting point to improve on existing heuristics for LDES bidding rules [brown_price_2025, dupre_la_tour_towards_2023]. They should be built on the insight that conditional marginal storage value derives from the cost and probability of extreme system states, which in turn varies with the seasonal underlying weather distribution.
@@
The introduction of limited foresight has also important implications for price formation. Not least the energy crisis in the wake of the Russian invasion of Ukraine has reignited a debate on electricity market design. Price formation on electricity markets with very high shares of variable renewables has been called into question as a substantial price disparity with long periods of near-zero prices and sporadic extreme prices threatens market functioning, cost recovery and raises financing costs [mallapragada_electricity_2022]. The prices in the perfect foresight model above resemble these patterns. However, once we introduce limited foresight and the long-duration storage bids a marginal storage value that represents the probability-adjusted costs of extreme system states, we observe much smoother price duration curves that may solve many of the aforementioned concerns. These findings are similar to those of [aaslid_pricing_2021] in a hydroelectric context and complement recent contributions by Brown et al. [brown_price_2025] and Antweiler and Muesgens [antweiler_new_2025] who highlight the role of demand elasticity and storage.
@@ 5.3 Limitations
The presented analysis is aimed to illustrate the mechanism through which weather uncertainty and LDES storage operations affect renewable energy systems as well as to give a sense on the order of magnitude of the effects. Nevertheless, there are several limitations that need to be considered when interpreting the results.
First, the case study adopts a very simplified energy system representation to maintain computational tractability. Each model is confined to a single country-node. Assuming away all interconnection with other countries in the European network is likely to exaggerate LDES requirements of extreme system states ([roth_geographical_2023, kittel_coping_2025]. Also the representation of sector coupling is overly simplified. For example, long-duration thermal storage in the heating sector may provide a significant amount of the flexibility provided by LDES in the model [schmidt_european_2025]. It has been shown that sector coupling plays a smoothing role in price formation, which may mitigate the extreme price bifurcation seen in the perfect foresight model above [brown_price_2025]. Likewise, we omit flexible hydro capacities.
@@
Third, we assume monthly stage lengths to strike a balance between the stagewise independence assumption and a realistic degree of foresight. The latter is likely to be overly optimistic. However, overestimating foresight probably results in a lower bound to the differences between perfect and limited foresight and we would expect effects to be stronger over shorter foresight horizons. A violation of the stagewise independence assumption leads to meteorologically implausible weather years that lead to substantial biases.
Fourth, the results indicate that there is a significant end-of-horizon effect ([dowson_policy_2020]). The LDES does not see any value in adding storage levels beyond the target level at 1st July. In a multi-year or infinite-horizon formulation the next winter would induce the storage to start stockpiling again earlier. The seasonal differences between bidding curves would likely be less pronounced as bids come down for autumn months and come up for spring and early summer months.
@@ 5.4 Future research
This analysis is meant to give an insight on the consequences of LDES bidding under weather uncertainty. We see several avenues for future research:
First, one could further decompose the role of different marginal weather distributions. For example, one could try to isolate the effect of heat demand uncertainty on LDES bidding curves. Furthermore, other research has found significant trade-offs between long-duration electricity storage and long-duration thermal storage [schmidt_european_2025]. A consideration of weather uncertainty promises further insights as it accounts for optionality. Second, the LDES bidding curves trained on historical wind and solar data could serve as a starting point to build more refined heuristic bidding rules like those suggested by Brown et al. and Dupre la Tour [brown_price_2025, dupre_la_tour_towards_2023], to emulate limited LDES foresight in large-scale capacity expansion models. Resulting non-linearities could be addressed with piece-wise linear approximations that have been frequently used in the context of technological learning [behrens_reviewing_2024]. Another avenue could be to use linear decision rules to approximate multi-stage decision-making [egging_linear_2017, nazare_solving_2023]. Third, the role of LDES bidding under uncertainty for price formation in fully renewable systems should be investigated further. There are likely to be interactions with demand elasticity, sector-coupled demand-side flexibility such as long-duration thermal storage in district heating networks and interconnection and may have consequences for optimal market design in a world with only VRE and storages [billimoria_contract_2023].
@@
The open-source code used in this analysis is available on Github at https://github.com/FSchmidtDIW/DIETER-SDDP. All results data are available in a Zenodo repository at (to be added). All input data can be fully reproduced using the open-source repo DieterData available in a Gitlab repo.
@@ Acknowledgments
This section provides a formal description of the multi-stage stochastic program that represents the Limited Foresight model. For simplicity we focus on the electricity sector and abstract the heat balance and simply model include electricity demand from heat in the electricity energy balance abstracting from some short-term flexibility in the heating sector included in the case study model. We also do not show the hydrogen balance in which hydrogen supply enters by way of electrolysis or imports and hydrogen demand arises from exogenous industrial demand and hydrogen gas turbines. Instead we present the long-duration energy storage consisting of electrolysis, caverns and H2 turbines as if it were a generic electricity storage technology and abstract from its integration in the hydrogen sector. Again, the case study of Section 3 is more detailed and represents the hydrogen sector explicitly, also accounting for electricity demand from compressors as described in [kittel_coping_2025].
The problem has $T+1$ stages, where stage 00 is the capacity investment stage and $t=1,\dots,T$ are dispatch dispatch stages. In the case study these are assumed to cover a month of four-hourly periods each resulting in stage lengths of $H(t)=|\mathcal{H}(t)|$ between 168 and 186 time steps.
@@
$\min_{\{G_{r},F_{s},H_{s},E_{s},e_{s}^{ini}\}\in\mathcal{C}}\sum_{r}c_{r}G_{r}+\sum_{s}\left(c_{s}^{f}F_{s}+c_{s}^{h}H_{s}+c_{s}^{e}E_{s}\right)+p_{LTC}I_{H2}$ (SI.1a)
$+\mathbb{E}_{0}\Bigg{[}\min_{\begin{subarray}{c}\{g_{r,1,h},f_{s,1,h},h_{s,1,h},g^{inf}_{1,h},e_{s,1,h}\}\\ \in\mathcal{O}_{1}(G_{r},F_{s},H_{s},E_{s},\xi_{1})\end{subarray}}\sum_{h\in\mathcal{H}(1)}\sum_{r}o_{r}g_{r,1,h}+vg^{inf}_{1,h}+\cdots$
$+\mathbb{E}_{T-1}\bigg{[}\min_{\begin{subarray}{c}\{g_{r,T,h},f_{s,T,h},h_{s,T,h},g^{inf}_{T,h},e_{s,T,h}\}\\ \in\mathcal{O}_{T}(G_{r},F_{s},H_{s},E_{s},e_{s,T-1},\xi_{T})\end{subarray}}\sum_{h\in\mathcal{H}(T)}\sum_{r}o_{r}g_{r,T,h}+vg^{inf}_{T,h}+\sum_{s}c^{terminal}(e_{s,T,H})\bigg{]}\Bigg{]}$
$c_{r}$ are the capital costs for generator capacity of type $r$, $c^{f}_{s},c^{h}_{s}$ and $c^{e}_{s}$ are the capital costs for dispatch power, charging power and energy capacity of storage type $s$ respectively. $G_{r}$ is the generation capacity choice, $F_{s},H_{s}$ and $E_{s}$ are the power and energy capacity choices for storage $s$. $\mathcal{C}$ is the compact and convex feasible set for capacity choices given by,
$\mathcal{C}:=\Bigg{\{}\{G_{r},F_{s},H_{s},E_{s}\}\Bigg{|}\bigg{\{}\begin{subarray}{c}0\leq G_{r}\leq\bar{G_{r}}\\ 0\leq F_{s}\leq\bar{F_{s}}\\ 0\leq H_{s}\leq\bar{H_{s}}\\ 0\leq E_{s}\leq\bar{E_{s}}\end{subarray}\bigg{\}}\forall r,s\Bigg{\}}.$ (SI.1b)
Note that a subset of storage $\mathcal{S}^{LDES}\subset\mathcal{S}$ are long-duration energy storages. The remainder are short-duration storage technologies for which the model does not introduce state variables. We assume that $\mathcal{S}^{LDES}$ is a singleton, consisting of PEM electrolysis, cavern storage and H2 turbines, but an inclusion of e.g. hydro reservoirs would also be possible. The dispatch stages $t=1,\dots,T$ are random and enter Eq. (SI.1) in expectation over the stochastic weather process $\{\xi_{t}\}_{t=1}^{T}$. The random weather vector $\xi_{t}$ affects the feasible space at dispatch stage $t$, $\mathcal{O}_{t}$. $h\in\mathcal{H}(t)$ is a time step in dispatch stage $t$. $\xi_{t}$ contains capacity factors for wind, onshore and offshore, and solar PV, $\phi_{r,t,h}$, as well as heat demand $d^{heat}_{t,h}$ and heat pump efficiencies $\psi^{hp}_{t,h}$. $g_{r,t,h}$ is the generation of generator $r$ in dispatch stage $t$ in period $h$. $f_{s,t,h},h_{s,t,h},e_{s,t,h}$ are the corresponding storage dispatch, charging and energy level choices for storage type $s$. $o_{r}$ are the marginal costs of generator $r$. $g^{inf}_{t,h}$ is a loss of load variable that is priced at value of lost load (VOLL), $v$. The feasible set of dispatch stage $t$ is given by,
@@
$d_{t,h}+\psi^{hp}_{t,h}d_{t,h}^{heat}+\sum_{s}h_{s,t,h}-\sum_{r}g_{r,t,h}-\sum_{s}f_{s,t,h}=0$ $\perp\lambda_{t,h}\in\mathbb{R}\quad\forall h\in\mathcal{H}(t)$ (SI.1d)
@@
[…]
The stages are connected by the capacity choices in stage $t=0$, $G_{r},H_{s},F_{s},E_{s}$ and the initial storage level $e_{s}^{ini}$ as well as the plant equation Eq. (SI.1o) propagating the storage level through the stages. The plant equation only needs to hold for storages $s\in\mathcal{S}^{LDES}$, the subset of storages that are LDES. The remaining storages are subject to a circularity constraint within each dispatch stage that is omitted here for simplicity. The greek symbols following the perpendicular sign, $\perp$, are the dual variables to the constraints. We will use them in the derivation of the storage bidding curve in Section 2.4.
@@
The Perfect Foresight Model that serves as a benchmark in isolating the effect of limited foresight under weather uncertainty is a two-stage stochastic program and a special case of the multi-stage stochastic program discussed in the previous subsection. The problem stages are $t=0,1$, where stage 00 is the capacity stage and stage $1$ is the dispatch stage. Stage $1$ has 2190 four-hourly time steps to cover an entire weather year. The number of scenarios is $N=N_{1}$, equal to 35 in the case study. Capacity decisions and the choice of the initial storage level are taken under uncertainty of which weather year is going to materialize in the dispatch stage. Otherwise the system is operated under perfect foresight.
@@
This section provides a very brief introduction to the stochastic dual dynamic programming (SDDP) algorithm used to solve the Problem in Eq. SI.1 for the Limited Foresight model in the main body of the text. The section builds on Dowson (2020) [dowson_policy_2020] and Shapiro (2011) [shapiro_analysis_2011] as well as Shapiro et al. (2021) [shapiro_lectures_2021] and we refer to these texts for a more thorough treatment.
@@
$\min_{\begin{subarray}{c}g_{r,t,h},f_{s,t,h},h_{s,t,h},g^{inf}_{t,h},e_{s,t,h}\\ \in\mathcal{O}_{t}(G_{r},F_{s},H_{s},E_{s},e_{s,t-1},\xi_{t})\end{subarray}}\sum_{h}\left[\sum_{r}o_{r}g_{r,t,h}+vg^{inf}_{t,h}\right]$ $+\mathbb{E}_{t}\left[{Q}_{t+1}(e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini})\right]$
@@
    Relatively complete recourse: We assume that regardless of decisions in stages $11,\dots,t-$ 1 there exists a feasible solution in stage $t$, i.e. $\mathcal{O}_{t}\neq\emptyset$
@@
Assuming that each stage has a number of samples $N_{t}$ the total number of scenarios given the stagewise independence assumption is $N=\prod_{t=1}^{T}$. Based on assumption 3, we replace the expectation operator in SI.2 by $\mathcal{Q}_{t}(\cdot)=\frac{1}{N_{t}}\sum_{i}^{N_{t}}Q_{t}(\cdot,\xi_{t,i}),t=1,\dots,T$.
By linearity of the subproblems, the cost-to-go functions $\mathcal{Q}_{t}(\cdot)$ are convex, piecewise-linear, in the state variables. We decompose the model by replacing the expected cost-to-go in each stage $t$ by a lower approximation of the same, a cutting-plane model. For simplicity lets subsume all outgoing state variables in a vector-valued variable $x_{t}=\{e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini}\}$ and all dispatch variables as $y_{t}=\{g_{r,t,h},f_{s,t,h},h_{s,t,h},g^{inf}_{t,h},e_{s,t,h}\}$. Given an arbitrary iteration $k$ of the algorithm, a weather realization $\xi_{t,i}$ and an incoming state $\bar{x}_{t-1}$, we can write the subproblem in stage $t$ as,
@@
Note that we have replaced the cost-to-go term by $\theta_{t}$, an approximation. We also simplify the representation of the current stage’s dispatch costs by $c_{t}^{T}y_{t}$. The cost vector $c_{t}$ only has non-zero entries for dispatchable generation with marginal generation costs and for load-shedding (or imports in the case study). Eq. SI.3b defines the feasible space for dispatch decisions $y_{t}$ given the weather realization $\xi_{t,i}$ and incoming state variable vector $x_{t-1}$. We take the detour of defining a fishing constraint Eq.SI.3c to conveniently collect the vector of duals with respect to the incoming state variables, which we define as $\rho_{t}$ [dowson_policy_2020]. Eq. SI.3d describes the transition of state variables. For capacities and the initial storage level $e_{s}^{ini}$ functional $T(\cdot)$ just passes on without altering anything. The transition from $e_{s,t-1,H}$ to $e_{s,t,H}$ is defined by Eq. SI.1o and SI.1e. Eq. SI.3e describes the cutting plane model. $\theta_{t}$ is bounded from below by a collection of cutting planes or cuts, defined by subgradients of $\mathcal{Q}_{t+1}$ obtained in previous iterations $j=1,\dots,k$ [^5] Since the cost-to-go functions are piecewise-linear they are not differentiable and we require the generalization of a gradient, that is a subgradient. See [goke_stabilized_2024] for details. The last constraint just ensures that our cost estimate is non-negative.
The more cuts we add to the approximate subproblem the closer $\theta_{t}$ resembles $\mathcal{Q}_{t+1}(\cdot)$ and the approximation converges to the true cost-to-go function. Next, we discuss how to add new cuts in an iteration. Each iteration consists of two steps, a forward pass and a backward pass.
@@
We need to create trial points $\{x_{t}^{k}\}_{t=0}^{T}$. In iteration $k$, we create a random sample path $\xi_{i}$ and begin by solving the approximate subproblem at stage 00. We pass on $x^{k}_{0}$ to stage $1$, observe realization $\xi_{1,i}$ and solve the approximate subproblem in Eq. SI.3, pass on $x^{k}_{1}$ to stage $2$, observe $\xi_{2,i}$ and so on. Note that the relatively complete recourse assumption ensures that the next stage’s subproblem remains feasible regardless of the incoming state variable or the weather realization.

### v2

@@ 5.2 Implications
There is a broad literature consensus that the design of variable renewable energy systems requires taking into account interannual weather variability [pfenninger_dealing_2017, collins_impacts_2018, ruggles_planning_2024, gotske_designing_2024, goke_stabilized_2024, grochowicz_intersecting_2023, grochowicz_using_2024], especially in the context of long-duration energy storage (LDES) serving as backup during prolonged periods of energy scarcity [brown_ultra-long-duration_2023, dowling_role_2020]. However, the role of limited foresight with respect to weather uncertainty in capacity expansion models is often ignored. Few studies have considered limited foresight in LDES operations and have either done so by using heuristic bidding rules [brown_price_2025, dupre_la_tour_towards_2023], or by using very reduced scenario trees [blanchard_strategic_2024, hummelen_exploring_2024] in partial capacity expansion models.
@@
Ignoring weather uncertainty, capacity expansion planning can lead to biased systems designs and operations. Perfect foresight models tend to rely too much on wind capacities, as they can perfectly leverage intermittent windy episodes throughout an operational year. In a limited foresight model, solar PV gains in system value as its capacity factors are less variable, aiding the LDES stockpiling process. Meanwhile, LDES capacities themselves do not change much, and energy capacities remain close to those defined by the largest dispatch events such as the winters of 1996/97 and 2010/11 under PF as found in previous literature contributions [kittel_coping_2025, ruhnau_storage_2022].
@@
It is obvious that highly resolved sector-coupled capacity expansion models would be intractable to solve with a stochastic (dual) dynamic programming method like we apply here. Yet, the derived LDES bidding curves may serve as a starting point to improve on existing heuristics for LDES bidding rules [brown_price_2025, dupre_la_tour_towards_2023]. They should be built on the insight that conditional marginal storage value derives from the cost and probability of extreme system states, which in turn varies with the underlying seasonal weather distribution.
@@
The introduction of limited foresight also has important implications for price formation. The energy crisis in the wake of the Russian invasion of Ukraine has reignited a debate on electricity market design. Price formation on electricity markets with very high shares of variable renewable energy has been called into question as a substantial price disparity with long periods of near-zero prices and sporadic extreme prices threatens market functioning, cost recovery and raises financing costs [mallapragada_electricity_2022]. The prices in the perfect foresight model above resemble these patterns. However, once we introduce limited foresight and the long-duration storage bids a marginal storage value that represents the probability-adjusted costs of extreme system states, we observe much smoother price duration curves that may solve many of the aforementioned concerns. These findings are similar to those of [aaslid_pricing_2021] in a hydroelectric context and complement recent contributions by Brown et al. [brown_price_2025] and Antweiler and Muesgens [antweiler_new_2025] who highlight the role of demand elasticity and storage.
@@ 5.3 Limitations
The presented analysis is aimed at illustrating the mechanism through which weather uncertainty and LDES storage operations affect renewable energy systems and giving a sense of the order of magnitude of the effects. Nevertheless, there are several limitations that need to be considered when interpreting the results.
First, the case study adopts a very simplified energy system representation to maintain computational tractability. Each model is confined to a single country-node. Assuming away all interconnection with other countries in the European network is likely to exaggerate LDES requirements of extreme system states [roth_geographical_2023, kittel_coping_2025]. Also the representation of sector-coupling is overly simplified. For example, long-duration thermal storage in the heating sector may provide a significant amount of the flexibility provided by LDES in the model [schmidt_mix_2025]. It has been shown that sector-coupling plays a smoothing role in price formation, which may mitigate the extreme price bifurcation seen in the perfect foresight model above [brown_price_2025]. Likewise, we omit flexible hydro capacities.
@@
Third, we assume monthly stage lengths to strike a balance between the validity of the stagewise independence assumption and a realistic degree of foresight. The latter is likely to be overly optimistic. However, overestimating foresight probably results in a lower bound to the differences between perfect and limited foresight and we would expect effects to be stronger over shorter foresight horizons. A violation of the stagewise independence assumption leads to meteorologically implausible weather years that may lead to substantial biases.
Fourth, the results indicate that there is a significant end-of-horizon effect [dowson_policy_2020]. The LDES does not see any value in adding storage levels beyond the target level at 1st July. In a multi-year or infinite-horizon formulation, the next winter would induce the storage to start stockpiling again earlier. The seasonal differences between bidding curves would likely be less pronounced as bids come down for autumn months and come up for spring and early summer months.
@@ 5.4 Future research
This analysis is meant to give an insight into the consequences of LDES bidding under weather uncertainty. We see several avenues for future research:
First, one could further investigate the role of different marginal weather distributions. For example, one could seek to isolate the effect of heat demand uncertainty on LDES bidding curves. Second, the LDES bidding curves trained on historical wind and solar data could serve as a starting point to build more refined heuristic bidding rules like those suggested by Brown et al. and Dupre la Tour [brown_price_2025, dupre_la_tour_towards_2023], to emulate limited LDES foresight in large-scale capacity expansion models. Resulting non-linearities could be addressed with piece-wise linear approximations that have been frequently used in the context of technological learning [behrens_reviewing_2024]. Another avenue could be to use linear decision rules to approximate multi-stage decision-making [egging_linear_2017, nazare_solving_2023]. Third, the role of LDES bidding under uncertainty for price formation in fully renewable systems should be investigated further. There are likely to be interactions with demand elasticity, sector-coupled demand-side flexibility, such as long-duration thermal storage in district heating networks [schmidt_mix_2025], and interconnection. It may have consequences for optimal market design in a system with only VRE and storage [billimoria_contract_2023].
@@
The open-source code used in this analysis is available on Github at https://github.com/FSchmidtDIW/DIETER-SDDP. All results data are available in a Zenodo repository at https://zenodo.org/records/15570551. All input data can be fully reproduced using the open-source repo DieterData available in a https://gitlab.com/diw-evu/dieter_public/dieterdata.
@@ Acknowledgments
This section provides a formal description of the multi-stage stochastic program that represents the Limited Foresight model. For simplicity, we focus on the electricity sector and abstract from the heat balance and simply include electricity demand from heat in the electricity energy balance. That way, we also abstract from some short-term flexibility in the heating sector included in the case study model. We also do not show the hydrogen balance in which hydrogen supply enters by way of electrolysis or imports and hydrogen demand arises from exogenous industrial demand and hydrogen gas turbines. Instead, we present the long-duration energy storage consisting of electrolysis, caverns and H2 turbines as if it were a generic electricity storage technology and abstract from its integration in the hydrogen sector. Again, the case study of Section 3 is more detailed and represents the hydrogen sector explicitly, also accounting for electricity demand from compressors as described in Kittel et al. [kittel_coping_2025].
The problem has $T+1$ stages, where stage 00 is the capacity investment stage and $t=1,\dots,T$ are dispatch stages. In the case study, these are assumed to cover a month of four-hourly periods, each resulting in stage lengths of $H(t)=|\mathcal{H}(t)|$ between 168 and 186 time steps.
@@
$\min_{\{G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},I_{H2}\}\in\mathcal{C}}\sum_{r}c_{r}G_{r}+\sum_{s}\left(c_{s}^{f}F_{s}+c_{s}^{h}H_{s}+c_{s}^{e}E_{s}\right)+p_{LTC}I_{H2}$ (SI.1a)
$+\mathbb{E}_{0}\Bigg{[}\min_{\begin{subarray}{c}\{g_{r,1,h},f_{s,1,h},h_{s,1,h},g^{inf}_{1,h},e_{s,1,h}\}\\ \in\mathcal{O}_{1}(G_{r},F_{s},H_{s},E_{s},\xi_{1})\end{subarray}}\sum_{h\in\mathcal{H}(1)}\left[\sum_{r}o_{r}g_{r,1,h}+vg^{inf}_{1,h}\right]+\mathbb{E}_{1}\bigg{[}\cdots$
$+\mathbb{E}_{T-1}\bigg{[}\min_{\begin{subarray}{c}\{g_{r,T,h},f_{s,T,h},h_{s,T,h},g^{inf}_{T,h},e_{s,T,h}\}\\ \in\mathcal{O}_{T}(G_{r},F_{s},H_{s},E_{s},e_{s,T-1},\xi_{T})\end{subarray}}\sum_{h\in\mathcal{H}(T)}\left[\sum_{r}o_{r}g_{r,T,h}+vg^{inf}_{T,h}\right]+\sum_{s}c_{s}^{terminal}\bigg{]}\bigg{]}\Bigg{]}$
$c_{r}$ are the capital costs for generator capacity of type $r$, $c^{f}_{s},c^{h}_{s}$ and $c^{e}_{s}$ are the capital costs for dispatch power, charging power and energy capacity of storage type $s$ respectively. $p_{LTC}$ is the long-term contract price for hydrogen imports. $G_{r}$ is the generation capacity choice, $F_{s},H_{s}$ and $E_{s}$ are the power and energy capacity choices for storage $s$. $I_{H2}$ is the selected long-term contract volume. $\mathcal{C}$ is the compact and convex feasible set for capacity choices given by,
$\mathcal{C}:=\left\{\{G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},I_{H2}\}\Bigg{|}\left\{\begin{subarray}{c}0\leq G_{r}\leq\bar{G_{r}}\\ 0\leq F_{s}\leq\bar{F_{s}}\\ 0\leq H_{s}\leq\bar{H_{s}}\\ 0\leq E_{s}\leq\bar{E_{s}}\\ 0\leq e^{ini}_{s}\leq E_{s}\\ 0\leq I_{H2}\leq\bar{I}\end{subarray}\right\}\forall r,s\right\}.$ (SI.1b)
Note that a subset of storage $\mathcal{S}^{LDES}\subset\mathcal{S}$ are long-duration energy storages. The remainder are short-duration storage technologies for which the model does not introduce state variables. In this simplified formalization, we assume that $\mathcal{S}^{LDES}$ is a singleton, consisting of PEM electrolysis, cavern storage and H2 turbines, but an inclusion of e.g. hydro reservoirs would also be possible. The case study above also includes the option of more expensive tank storage. The dispatch stages $t=1,\dots,T$ are random and enter Eq. SI.1 in expectation over the stochastic weather process $\{\xi_{t}\}_{t=1}^{T}$. The random weather vector $\xi_{t}$ affects the feasible space at dispatch stage $t$, $\mathcal{O}_{t}$. $h\in\mathcal{H}(t)$ is a time step in dispatch stage $t$. $\xi_{t}$ contains capacity factors for wind, onshore and offshore, and solar PV, $\phi_{r,t,h}$, as well as heat demand $d^{heat}_{t,h}$ and heat pump efficiencies $\psi^{hp}_{t,h}$. $g_{r,t,h}$ is the generation of generator $r$ in dispatch stage $t$ in period $h$. $f_{s,t,h},h_{s,t,h},e_{s,t,h}$ are the corresponding storage dispatch, charging and energy level choices for storage type $s$. $o_{r}$ are the marginal costs of generator $r$. $g^{inf}_{t,h}$ is a loss of load variable that is priced at the value of lost load (VOLL), $v$. The feasible set of dispatch stage $t$ is given by,
@@
$d_{t,h}+(\psi^{hp}_{t,h})^{-1}d_{t,h}^{heat}+\sum_{s}h_{s,t,h}-\sum_{r}g_{r,t,h}-\sum_{s}f_{s,t,h}=0$ $\perp\lambda_{t,h}\in\mathbb{R}\quad\forall h\in\mathcal{H}(t)$ (SI.1d)
@@
Note that the weather uncertainty enters each dispatch stage in Eq. SI.1d for heat demand and in Eq. SI.1g for renewable capacity factors. We assume that $\phi_{r,t,h}$ are constant for $r\in\{bio,nuclear\}.$
The stages are connected by the capacity choices in stage $t=0$, $G_{r},H_{s},F_{s}$ and $E_{s}$ and the initial storage level $e_{s}^{ini}$ as well as the plant equation Eq. SI.1o propagating the storage level through the stages. The plant equation only needs to hold for storages $s\in\mathcal{S}^{LDES}$. The remaining storage technologies are subject to a circularity constraint within each dispatch stage that is omitted here for simplicity. The Greek symbols following the perpendicular sign, $\perp$, are the dual variables to the constraints. We use them in the derivation of the storage bidding curve in Section 2.4.
@@
The Perfect Foresight Model which serves as a benchmark in isolating the effect of limited foresight under weather uncertainty, is a two-stage stochastic program and a special case of the multi-stage stochastic program discussed in the previous subsection. The problem stages are $t=0,1$, where stage 00 is the capacity stage and stage $1$ is the dispatch stage. Stage $1$ has 2190 four-hourly time steps to cover an entire weather year. The number of scenarios is $N=N_{1}$, equal to 35 in the case study. Capacity decisions and the choice of the initial storage level are taken under uncertainty of which weather year is going to materialize in the dispatch stage. Otherwise, the system is operated under perfect foresight.
@@
This section provides a very brief introduction to the stochastic dual dynamic programming (SDDP) algorithm used to solve the Problem in Eq. SI.1 for the Limited Foresight model in the main body of the text. The section builds on Dowson (2020) [dowson_policy_2020] and Shapiro (2011) [shapiro_analysis_2011] as well as Shapiro et al. (2021) [shapiro_lectures_2021], and we refer to these texts for a more thorough treatment.
@@
$\min_{\begin{subarray}{c}g_{r,t,h},f_{s,t,h},h_{s,t,h},g^{inf}_{t,h},e_{s,t,h}\\ \in\mathcal{O}_{t}(G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},e_{s,t-1},\xi_{t})\end{subarray}}\sum_{h}\left[\sum_{r}o_{r}g_{r,t,h}+vg^{inf}_{t,h}\right]$ $+\mathbb{E}_{t}\left[{Q}_{t+1}(e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini})\right]$
@@
    Relatively complete recourse: We assume that regardless of decisions in stages $11,\dots,t-$ 1 and of the realization of $\xi_{t}$, there exists a feasible solution in stage $t$, i.e. $\mathcal{O}_{t}\neq\emptyset$
@@
Assuming that each stage has a number of samples $N_{t}$, the total number of scenarios given the stagewise independence assumption is $N=\prod_{t=1}^{T}N_{t}$. Based on assumption 3, we replace the expectation operator in SI.2 by $\mathcal{Q}_{t}(\cdot)=\frac{1}{N_{t}}\sum_{i}^{N_{t}}Q_{t}(\cdot,\xi_{t,i}),t=1,\dots,T$.
By linearity of the subproblems, the cost-to-go functions $\mathcal{Q}_{t}(\cdot)$ are convex, piecewise-linear, in the state variables. We decompose the model by replacing the expected cost-to-go in each stage $t$ by a lower approximation of the same, a cutting-plane model. For simplicity lets subsume all outgoing state variables in a vector-valued variable $x_{t}=\{e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini}\}$ and all dispatch variables as $y_{t}=\{g_{r,t,h},f_{s,t,h},h_{s,t,h},g^{inf}_{t,h},e_{s,t,h}\}$. Given an arbitrary iteration $k$ of the algorithm, a weather realization $\xi_{t,i}$ and an incoming state $x_{t-1}$, we can write the subproblem in stage $t$ as,
@@
Note that we have replaced the cost-to-go term by $\theta_{t}$, an approximation. We also simplify the representation of the current stage’s dispatch costs by $c_{t}^{T}y_{t}$. The cost vector $c_{t}$ only has non-zero entries for dispatchable generation with marginal generation costs and for load-shedding (or imports in the case study). Eq. SI.3b defines the feasible space for dispatch decisions $y_{t}$ given the weather realization $\xi_{t,i}$ and incoming state variable vector $x_{t-1}$. We define a fishing constraint Eq. SI.3c to conveniently collect the vector of duals with respect to the incoming state variables, which we define as $\rho_{t}$ [dowson_policy_2020]. Eq. SI.3d describes the transition of state variables. For capacities and the initial storage level $e_{s}^{ini}$ functional $T(\cdot)$ just passes on without altering anything. The transition from $e_{s,t-1,H}$ to $e_{s,t,H}$ is defined by Eq. SI.1o and SI.1e. Eq. SI.3e describes the cutting plane model. $\theta_{t}$ is bounded from below by a collection of cutting planes or cuts, defined by subgradients of $\mathcal{Q}_{t+1}$ obtained in previous iterations $j=1,\dots,k$ [^5] Since the cost-to-go functions are piecewise-linear they are not differentiable and we require the generalization of a gradient, that is a subgradient. See [goke_stabilized_2024] for details. The last constraint just ensures that our cost estimate is non-negative.
The more cuts we add to the approximate subproblem, the closer $\theta_{t}$ resembles $\mathcal{Q}_{t+1}(\cdot)$ and the approximation converges to the true cost-to-go function. Next, we discuss how to add new cuts in an iteration. Each iteration consists of two steps, a forward pass and a backward pass.
@@
We need to create trial points $\{x_{t}^{k}\}_{t=0}^{T}$. At the start of iteration $k$, we create a random sample path $\xi_{i}$ and begin by solving the approximate subproblem at stage 00. We pass on $x^{k}_{0}$ to stage $1$, observe realization $\xi_{1,i}$ and solve the approximate subproblem in Eq. SI.3, pass on $x^{k}_{1}$ to stage $2$, observe $\xi_{2,i}$ and so on. Note that the relatively complete recourse assumption ensures that the next stage’s subproblem remains feasible regardless of the incoming state variable or the weather realization.